### v2

@@ 6.2 Multiple-lane roads
Figure 17: (a). Two independent lanes  (b). Two-lane road allowing lane-changing
In particular, we look into the scenario where two collisions occur on two adjacent lanes on a two-lane roadway. Figure 17 compares the movements of AV strings under cases when AVs are allowed to change lanes or not. If lane changes are allowed, a throughput equals the full capacity of one lane can still be maintained. Accordingly, the lane-average expected throughput is given by
@@
$\approx\frac{H^{2}}{\tau^{2}}\frac{L^{2}}{\eta^{2}v^{2}}p(v,\eta)^{2}$ (24b)
@@
$\approx\left(1-\frac{TL}{\tau\eta v}p(v,\eta)+\left(\left(HT-\frac{H^{2}}{2}+\frac{L(v+c)(H-T)}{12vc}\right)\frac{L^{2}-L\tau\eta v}{\tau^{2}\eta^{2}v^{2}}+\frac{T^{2}L^{2}}{2\tau^{2}\eta^{2}v^{2}}\right)p(v,\eta)^{2}\right)s^{+}(v,\eta)$ (24d)
@@ 6.3 Comparison
In the end, we compare the hourly expected throughput in the extended scenarios with CIC under the baseline single-lane scenario with non-overlapping collisions. For the fair comparison purpose, we first transform CIC in Eq (12) to its one-hour approximation form by considering only one collision happens in one hour. This approximation is based on the consideration that two independent collisions will block the lane for the entire hour, as each of their TCTs exceeds half on hour:
@@
$=\left(1-\frac{TL}{\tau\eta v}p(v,\eta)\right)s^{+}(v,\eta)$ (25)
Figure 18: The comparison of the one-hour expected throughput under three conditions
@@ 7 Conclusion
In this paper, we analyzed the impact of microscopic robotic errors in autonomous vehicles on macroscopic traffic in terms of collisions and capacity in the car-following scenario. The systematic errors inherent in AV operations, ranging from perception to control, contribute to their stochastic deviation from the intended movement trajectory. These random movements serve as a source of collisions, undermining the safety and mobility performance of the fully autonomous traffic system.
A modified IDM model is adopted to describe AVs’ car-following behaviors under robotic errors, revealing that the resulting car-following distance follows a Gaussian distribution. This characteristic is further validated using real-world autonomous driving data from Waymo. The statistical nature also allows for an explicit formulation of the probability of rear-end collisions caused by uncertainties in the car-following headway. By incorporating the total clearance time required for a collision resolution, the dynamics of traffic capacity at each point are mathematically modeled as a function of speed and headway through a semi-Markov process. The expected value of capacity, referred to as collision-inclusive capacity, serves as a key metric for evaluating the combined mobility and safety performance of fully autonomous traffic.
@@
Our future work will extend the investigation into safety and mobility trade-offs in more AV-involved traffic operational scenarios, including those with complex traffic dynamics and behaviors or those under AV-HDV mixed traffic. Given that Gaussian assumptions have been widely used to model the uncertainty of complex self-driving environments (Cao et al., 2023), analytical models are expected to be applicable in these scenarios as well. Additionally, data-driven techniques, such as kernel estimation and learning-based methods, can also be employed to establish vehicle behavior models and describe driving scenarios, further enhancing the universality of this approach. Notably, randomness in human-driven vehicles has been investigated for both macroscopic traffic flow models (Jabari and Liu, 2012) and microscopic car-following models (Xu and Laval, 2020). In conjunction with our work, this existing body of research provides a solid foundation for understanding the compound stochasticity in mixed AV-HDV traffic.

### v3

@@ 6.2 Multiple-lane roads
Figure 17: Two collisions occur on: (a) two independent lanes without lane-changing; (b) a two-lane road with lane-changing allowed.
In particular, we look into the scenario where two collisions occur on two adjacent lanes on a two-lane roadway. Figure 17 compares the movements of AV strings under cases when AVs are allowed to change lanes or not. If lane changes are allowed, a throughput equals the full capacity of one lane can still be maintained. Accordingly, the lane-average expected throughput is given by:
@@
$\approx\frac{H^{2}}{\tau^{2}}\frac{L^{2}}{\eta^{2}v^{2}}p(v,\eta)^{2};$ (23b)
@@
$\approx\left(1-\frac{TL}{\tau\eta v}p(v,\eta)+\left(\left(HT-\frac{H^{2}}{2}+\frac{L(v+c)(H-T)}{12vc}\right)\frac{L^{2}-L\tau\eta v}{\tau^{2}\eta^{2}v^{2}}+\frac{T^{2}L^{2}}{2\tau^{2}\eta^{2}v^{2}}\right)p(v,\eta)^{2}\right)s^{+}(v,\eta).$ (23d)
@@ 6.3 Comparison
Finally, we compare the hourly expected throughput in the extended scenarios with CIC derived under the baseline single-lane setting with non-overlapping collisions. Yet, these two metrics are not directly comparable, as the former represents an hourly expectation, while the latter reflects a long-term average under stationary conditions.To ensure a fair comparison, we first transform CIC in Eq (12) into its one-hour approximation by considering the case in which at most one collision occurs within an hour. This approximation is motivated by the consideration that two independent collisions would block the lane for the entire hour as each of their TCTs exceeds half an hour. Mathematically, the resulting expression corresponds to the first-order Taylor expansion of Eq (12). Consistent with the derivations shown in Eqs (22f) and (23d), it neglects higher-order terms, specifically those of order $p(v,\eta)^{2}$ and beyond. The final mathematically expression is given by:
@@
$=\left(1-\frac{TL}{\tau\eta v}p(v,\eta)\right)s^{+}(v,\eta).$ (24)
@@
Figure 18: The comparison of the one-hour expected throughput (in blue curves) under three conditions with the same single-vehicle collision probability.
@@ 7 Conclusion
In this paper, we analyzed the impact of microscopic robotic uncertainties in autonomous vehicles on macroscopic traffic in terms of collisions and capacity in the car-following scenario. These systematic uncertainties rooted in sensors, onboard algorithms, and actuators of all mobile robots including AVs, which are unrelated to the external environment and driving intentions, contribute to their stochastic deviation from the intended movement trajectory. These random movements serve as a source of collisions, undermining the safety and mobility performance of the fully autonomous traffic system.
A modified IDM model is adopted to describe AVs’ car-following behaviors under robotic errors, including perception and control errors, revealing that the resulting car-following distance follows a Gaussian distribution. This characteristic is further validated using real-world autonomous driving data from Waymo. The statistical nature also allows for an explicit formulation of the probability of rear-end collisions caused by uncertainties in the car-following headway. By incorporating the total clearance time required for a collision resolution, the dynamics of traffic capacity at each point are mathematically modeled as a function of speed and headway through a semi-Markov process. The expected value of capacity, referred to as collision-inclusive capacity, serves as a key metric for evaluating the combined mobility and safety performance of fully autonomous traffic.
@@
Our future work will extend the investigation into safety and mobility trade-offs in more AV-involved traffic operational scenarios, including those with complex traffic dynamics, as well as mixed traffic flows involving both AVs and HDVs, or heterogeneous AVs from different manufacturers. As robotic uncertainty is widely recognized in complex self-driving scenarios (Cao et al., 2023), our analytical model grounded in probabilistic characterizations is expected to remain applicable under these conditions.
In scenarios involving both AVs and HVs, it it worth noting that the randomness associated with human-driven vehicles has been investigated in both macroscopic traffic flow models (Jabari and Liu, 2012) and microscopic car-following models (Xu and Laval, 2020). Building on this existing body of research, our microscopic analysis can be extended to incorporate the heterogeneous motion stochasticity for AVs and HDVs, contributing to a promising foundation for modeling compound uncertainty in mixed AV-HDV traffic.